OpenStudy (anonymous):

Two identical 1.50kg masses are pressed against opposite ends of a spring of force constant 1.85N/cm , compressing the spring by 20.0cm from its normal length. Find the maximum speed of each mass when it has moved free of the spring on a smooth, horizontal lab table. **Can someone help me with the set up of this problem? I keep getting the wrong answer and don't know what I am doing wrong.**

OpenStudy (anonymous):

try energy conservation \[\frac{ 1 }{ 2 } m v ^{2} + \frac{ 1 }{ 2 } m v ^{2} = \frac{ 1 }{ 2 } k x ^{2}\] speed is max when spring attend it's natural length

OpenStudy (anonymous):

your answer would be in cm /s

OpenStudy (anonymous):

Would x be ten or twenty do you think?

OpenStudy (anonymous):

potential energy of spring=(kx^2)/2 where x is compression or elongation in it's natural length so it would be 20 cm

OpenStudy (anonymous):

I get 15.7cm/s when I type the numbers in, but it still says that it is incorrect.

OpenStudy (anonymous):

convert every cm in m.

OpenStudy (anonymous):

1.57 m/s

OpenStudy (anonymous):

thanks, that was the problem, I was just converting at the end.

OpenStudy (anonymous):

you are welcome.